# Cascade Speculative Drafting for Even Faster LLM Inference

### v4

@@ Abstract
Introduced to enhance the efficiency of large language model (LLM) inference, speculative decoding operates by having a smaller model generate a draft. A larger target model then reviews this draft to align with its output, and any acceptance by the target model results in a reduction of the number of the target model runs, ultimately improving efficiency. However, the drafting process in speculative decoding includes slow autoregressive generation and allocates equal time to generating tokens, irrespective of their importance. These inefficiencies collectively contribute to the suboptimal performance of speculative decoding. To further improve LLM inference, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that incorporates two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade optimizes time allocation in drafting for improved efficiency. Combining both cascades, CS Drafting achieves up to an 81 percent additional speedup over speculative decoding in our experiments, while maintaining the same output distribution as the target model. Our code is publicly available at https://github.com/lfsszd/CS-Drafting.
Machine Learning, ICML
[…]
@@ 1 Introduction
The advent of Large Language Models (LLMs), like GPT-4 (OpenAI, 2023), has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but have also found widespread applications in user-interactive settings, such as chatbots and virtual assistants. However, these applications involve an extremely high number of users, up to hundreds of millions daily. To serve in real-time at this scale, a low-latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the service provider and the community. Consequently, optimizing the latency of LLMs has become a critical area of research.
@@
Unfortunately, the ever-growing size of LLMs significantly increases the latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one. An emerging solution, known as speculative decoding (Leviathan et al., 2023; Chen et al., 2023; Xia et al., 2023), shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter) autoregressively, and these tokens are then reviewed by a target model (which is larger and slower) in parallel. In one single run, the target model will accept any tokens aligned with its output and further generate one token. The drafting process in speculative decoding enables the target model to generate multiple tokens in a single run while maintaining its output distribution unchanged. With a properly sized draft model, speculative decoding achieves a speedup of 2 to 3 times, making it a potential method for solving high latency issues.
@@
Figure 1: The figure shows an example of Cascade Speculative Drafting with target model $\mathcal{M}_{t}$ and draft models $\mathcal{M}_{d_{1}}$, $\mathcal{M}_{d_{2}}$, and $\mathcal{M}_{d_{3}}$. The horizontal cascade involves using larger draft models to generate the earlier tokens and smaller models to generate the later tokens. The vertical cascade requires each model to review drafts from smaller models with the exception of the smallest model which is a statistical language model. As the horizontal cascade and vertical cascade are orthogonal to each other, CS Drafting combines both approaches for optimal efficiency.
@@
However, since draft models are typically required to generate multiple tokens in multiple steps, where each generation still involves inefficient autoregressive decoding, the performance of speculative decoding could be limited by the drafting latency. This inefficiency is also indicated by Leviathan et al. (2023)’s experiments, where it was observed that very small models (e.g., around two orders of magnitude smaller than the target model) are usually the best choice for drafting because their inference cost is lower compared to that of a larger draft model, despite the fact that larger draft models usually have higher-quality generation. This underscores that improving drafting efficiency is crucial for further enhancing the performance of speculative decoding.
@@
Figure 2: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on GSM8K. The draft model generates 30 tokens at each step. The results on MMLU are shown in Figure 3 of the Appendix.
@@
Inspired by the above observations, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, starting with the smallest being a statistical language model. Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model. In this design, the drafting of each neural model will be accelerated by drafting from a smaller model, avoiding the inefficiency of autoregressive generation from neural models. We refer to this tiered speculative decoding approach as the Vertical Cascade. In addition, we suggest the use of smaller, faster draft models for generating high-rejection tokens that are trailing in drafting generation, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS Drafting approach, as illustrated in Figure 1.
@@ 2 Preliminary
The core concept of speculative decoding (Leviathan et al., 2023) involves the utilization of a small draft model for sequential token generation with validation by a larger target model resulting in reduced latency. This design accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models.
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, $k$ be the number of draft tokens generated per step, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$. In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]\leq\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $norm(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$. At the end, the target model will generate one additional token following the accepted tokens. Such a design guarantees the output is the same as sampling autoregressively using the target model alone, as proven in Leviathan et al. (2023).
@@ 3.1 Vertical Cascade
Additionally, we incorporate lenience, a hyperparameter that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model (Leviathan et al., 2023). Lenience can be adopted during sampling or greedy decoding with speculative decoding. Let lenience $l\in[1,\infty)$. When sampling, the acceptance condition for token $x_{i}$ is transformed to $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$. If the acceptance condition is not satisfied, with a probability of $1-\frac{l\times\mathcal{M}_{t}(x)}{\mathcal{M}_{d}(x)}$, we reject $x_{i}$ and any following tokens. [^1] For simplicity, we assume the probability outputs are not standardized. We refer the readers to Leviathan et al. (2023) for the discussion on standardized sampling. When performing greedy decoding, the acceptance condition becomes deterministic and is simply either $argmax\mathcal{M}_{d}(x)[i]=argmax\mathcal{M}_{t}(x)[i]$ or $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$.
@@ 4 Analysis
We begin with some notions. Let $\mathcal{M}_{t}$ be the target model, $\mathcal{M}_{d}$ be the draft model, and $k$ be the number of draft tokens generated per step.
@@
Despite the simple setting of EWIF, Leviathan et al. (2023) have demonstrated that it aligns with the experimental results in most instances. Therefore, our analysis will concentrate on EWIF.
@@ 4.1 Vertical Cascade
We analyze EWIF of vertical cascade using generating functions, a well-studied topic in combinatorial mathematics (West, 2021). The properties of generating functions are useful in the recursion and evaluation process making the our final expression simple.
@@
$\phi_{\alpha^{\prime},k}(\alpha)<\alpha$ for any $01>\alpha>0,1>\alpha^{\prime}>0,k>$ 0, so if $c_{d_{2}}<<1$, the EWIF of $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$ is higher than EWIF of $\mathcal{M}_{d_{1}}$ alone.
@@
Therefore, with the statistical model having negligible cost (i.e., $c_{d_{2}}<<1$), it can almost always improve the efficiency of an SD system.
@@ 4.2 Horizontal Cascade
Using the information provided by Leviathan et al. (2023), we calculate a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial. The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
@@
Table 1: Simulated EWIF under the assumption that the acceptance distribution is a Bernoulli distribution. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. In the simulation, speculative sampling with horizontal cascade exceeded the performance of the vanilla speculative decoding on both CNN Dailymail (Nallapati et al., 2016) and WMT EnDe (Bojar et al., 2018) datasets.
@@ 5 Experiments
5.1 Evaluation Metric
Previous works on speculative decoding and related methods relied on walltime as an evaluation method. However, there are standardization and legitimacy concerns related to walltime. Leviathan et al. (2023) mentioned that cost coefficiency, the ratio between the time for a single run of draft model and target model will vary from system to system which can result in significant variation in wall time improvement across different systems, making comparisons across different research work difficult.
In addition, a recent analysis suggests GPU speed can vary for the same GPU model with one being 1.5x faster than another GPU of the same model (Sinha et al., 2022); therefore, theoretical walltime improvement up to 1.5x can be undermined by GPU variation, and methods without any improvement can be measured with an improvement by running on a different GPU. The lack of standardization and stability makes walltime unfit for measuring algorithmic improvement. To resolve this issue, we design a new metric to measure speculative decoding and its related algorithms which can reduce the impact of varying/unstable GPU performance.
Our proposed method, standardized walltime improvement (SWI), calculates the GPU times of the models, assuming that each run of a language model costs the same amount of time, an assumption made when inventing the speculative sampling algorithm (Chen et al., 2023). Given a speculative decoding system $S$ containing $k$ draft models $\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{k}}$, target model $\mathcal{M}_{t}$ solving a task $T$, and let $count(S,m,T)$ be the count of runs of model $m$ during the execution of task $T$ with system $S$, let $c_{i}$ be the expected cost coefficient being the time ratio between a single run of $\mathcal{M}_{i}$ and $\mathcal{M}_{t}$, and let $n$ be the total of tokens generated, the standardized walltime improvement is defined as
$\begin{split}SWI&(\mathcal{M}_{d_{1}},...,\mathcal{M}_{d_{k}},\mathcal{M}_{t},T)=\\ &\frac{n}{count(S,\mathcal{M}_{t},T)+\sum_{i=1}^{n}count(S,\mathcal{M}_{d_{i}},T)c_{i}},\end{split}$
which is simply the sum of GPU time by all models when $c_{i}$ is measured during a local experiment. The benefit of the $SWI$ evaluation is that $count(S,\mathcal{M}_{t},T)$ and $count(S,\mathcal{M}_{d_{i}},T)$ will be the consistent across different systems, so by using the same $c_{i}$ across different experiments the GPU performance variation will be eliminated; therefore, it provides a standardized, reproducible measurement.
During our experiment, we will use two sets of $\{c_{i}\}$, existing $\{c_{i}\}$ reported by Leviathan et al. (2023) and an heuristic $\{c_{i}\}$ being the ratio of number of tunable parameters between $\mathcal{M}_{d_{i}}$ and $\mathcal{M}_{t}$.
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (MS) $k_{11}$ $k_{12}$ $k_{22}$ $l$
GSM8K Autoregressive - 1 - - - -
GSM8K S Decoding base $3.38$ 10 - - -
GSM8K S Decoding small $3.06$ 11 - - -
GSM8K CS Drafting base, mag $3.70$ 10 - - -
GSM8K CS Drafting small, mag $3.19$ 11 - - -
GSM8K CS Drafting base, small, mag 3.88 8 13 1 3
MMLU Autoregressive - 1 - - - -
MMLU S Decoding base $3.97$ 13 - - -
MMLU S Decoding small $4.12$ 19 - - -
MMLU CS Drafting base, mag $4.56$ 13 - - -
MMLU CS Drafting small, mag $4.39$ 14 - - -
MMLU CS Drafting base, small, mag 4.88 5 19 1 5
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (PW) $k_{11}$ $k_{12}$ $k_{22}$ $l$
GSM8K Autoregressive - 1 - - - -
GSM8K S Decoding base $2.99$ 8 - - -
GSM8K S Decoding small $2.76$ 8 - - -
GSM8K CS Drafting base, mag $3.27$ 9 - - -
GSM8K CS Drafting small, mag $2.82$ 11 - - -
GSM8K CS Drafting base, small, mag 3.43 5 9 1 3
MMLU Autoregressive - 1 - - - -
MMLU S Decoding base $3.42$ 10 - - -
MMLU S Decoding small $3.51$ 11 - - -
MMLU CS Drafting base, mag $4.21$ 6 - - -
MMLU CS Drafting small, mag $3.99$ 13 - - -
MMLU CS Drafting base, small, mag 4.32 5 8 1 5
@@ 5.1 Evaluation Metric
Table 2: The experimental results on FLAN-T5. Speedup (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size). Speedup (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work (Leviathan et al., 2023). $k_{11}$, $k_{12}$, $k_{22}$, $l$ are the hyperparameters. $k_{11}$ and $k_{12}$ represent the step limitation target model and the draft models, $k_{22}$ is the step limitations between the first and second draft model, and $l$ is lenience as shown in algorithm 1. For speculative decoding, the $k_{11}$ is simply the $k$.
5.2 Setup
To ensure the generality of our findings, we perform experiments on both encoder-decoder and decoder-only models. For encoder-decoder models, we choose our target and draft models from the FLAN-T5 (Chung et al., 2022) family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters). We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models. For decoder-only models, we select LLAMA-2-chat-7B (Touvron et al., 2023) as the target model. Since there is no official release of a smaller model in LLAMA family, we use a pretrained 160M model with the same tokenizer as the reviewing draft model (Miao et al., 2024). In both cases, the Max-Gram algorithm is used as the generating draft model. To align our experiment with current common usage, we do not perform fine-tuning, and the generation is conducted in a zero-shot manner.
Since we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments (Leviathan et al., 2023), and to ensure our experiments are fully reproducible, we perform sampling at temperature 00, i.e., using greedy decoding by default.
Cascade Speculative Drafting Inference To reduce the number of hyperparameters to tune, we use MaG to generate 10 tokens at once, as it is rare for more than 10 tokens to be accepted. We do not use lenience when the reviewer is $\mathcal{M}_{t}$ to ensure the output distribution does not change. We also avoid lenience between MaG and its reviewer, since there is still a significant performance gap between MaG and a neural model.
With these constraints, we are left with at most four hyperparameters: $k_{11}$, $k_{12}$, $k_{22}$, and $l$. For the CS Drafting step where the target model is the reviewer, $k_{11}$ and $k_{12}$ are used. $k_{21}$ and $l$ are used in the step where $\mathcal{M}_{d_{1}}$ is the reviewer.
Baselines We consider both speculative decoding and vanilla autoregressive generation as our baselines. When performing speculative decoding with the target model as FLAN-T5-xxl, we opt for FLAN-T5-small and FLAN-T5-base as the draft models, as they perform better than FLAN-T5-large, according to Leviathan et al. (2023) and our preliminary experiments. When LLAMA-2-chat-7B is the target model, we use the 160M model as the draft model.
We test a wide range of $k$-values, from 2 to 30, to ensure that we find the best result for each model serving as our draft model to establish our baseline.
@@ 5.2 Setup
Dataset We chose two commonly used datasets for our experiments. For both datasets, we conducted experiments in a zero-shot chain-of-thought setup (Kojima et al., 2022; Wei et al., 2023):
@@
  • GSM8K (Cobbe et al., 2021) is a dataset comprising 8,500 high-quality, linguistically diverse, grade-school math word problems. It focuses on multi-step reasoning with problems that are typically solvable using basic arithmetic in 2-8 steps.
@@
  • MMLU (Hendrycks et al., 2021), or Massive Multitask Language Understanding, is a benchmark for testing how well large language models grasp knowledge. It encompasses 57 diverse subjects, ranging from elementary science to advanced law.

### v5

@@ Abstract
Introduced to enhance the efficiency of large language model (LLM) inference, speculative decoding operates by having a smaller model generate a draft. A larger target model then reviews this draft to align with its output, and any acceptance by the target model results in a reduction of the number of the target model runs, ultimately improving efficiency. However, the drafting process in speculative decoding includes slow autoregressive generation and allocates equal time to generating tokens, irrespective of their importance. These inefficiencies collectively contribute to the suboptimal performance of speculative decoding. To further improve LLM inference, we introduce Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that incorporates two types of cascades. The Vertical Cascade eliminates autoregressive generation from neural models, while the Horizontal Cascade optimizes time allocation in drafting for improved efficiency. Combining both cascades, CS Drafting achieves greater speedup compared to the baselines in our experiments, while preserving the same output distribution as the target model. [^1] Code is publicly available at https://github.com/lfsszd/CS-Drafting.
@@ 1 Introduction
The advent of Large Language Models (LLMs), like GPT-4 [17], has marked a significant milestone in the field of natural language processing (NLP). These models have not only excelled in various NLP tasks but have also found widespread applications in user-interactive settings, such as chatbots and virtual assistants. However, these applications involve an extremely high number of users, up to hundreds of millions daily. To serve in real-time at this scale, a low-latency system is not only cost-saving but also crucial for keeping the service running. In addition, the sheer scale of the service means that even a slight improvement in the latency of LLMs can greatly contribute to both the service provider and the community. Consequently, optimizing the latency of LLMs has become a critical area of research.
@@
Unfortunately, the ever-growing size of LLMs significantly increases the latency, especially in long-form generation, as autoregressive LLMs generate tokens one by one. An emerging solution, known as speculative decoding [14, 4, 25], shows potential to mitigate this issue. In speculative decoding, a draft model (which is smaller and faster) generates $k$ tokens in each step (with $k$ being a hyperparameter) autoregressively, and these tokens are then reviewed by a target model (which is larger and slower) in parallel. In one single run, the target model will accept any tokens aligned with its output and further generate one token. The drafting process in speculative decoding enables the target model to generate multiple tokens in a single run while maintaining its output distribution unchanged. With a properly sized draft model, speculative decoding achieves a speedup of 2 to 3 times, making it a potential method for solving high latency issues.
@@
Figure 1: The CS Drafting algorithm features a recursive and resource-efficient design, implemented through two cascades: the horizontal cascade and the vertical cascade. The horizontal cascade involves using larger draft models to generate the earlier tokens and smaller models for the later tokens. The vertical cascade requires each model to review drafts from smaller models with the exception of the smallest model, which is a statistical language model. As the horizontal cascade and vertical cascade are orthogonal, CS Drafting combines both approaches for optimal efficiency. The figure shows an example of Cascade Speculative Drafting with target model $\mathcal{M}_{t}$ and draft models $\mathcal{M}_{d_{1}}$, $\mathcal{M}_{d_{2}}$, and $\mathcal{M}_{d_{3}}$.
@@
However, since draft models are typically required to generate multiple tokens in multiple steps, where each generation still involves inefficient autoregressive decoding, the performance of speculative decoding could be limited by the drafting latency. This inefficiency is also indicated by Leviathan et al. [14], where it was observed that very small models (e.g., around two orders of magnitude smaller than the target model) are usually the best choice for drafting because their inference cost is lower compared to that of a larger draft model, despite the fact that larger draft models usually have higher-quality generation. This underscores that improving drafting efficiency is crucial for further enhancing the performance of speculative decoding.
@@
Figure 2: The probability of acceptance of draft tokens in relation to their positions in a single step of speculative decoding, evaluated on FLAN-T5-small, base, and large models on GSM8K and MMLU. The draft model generates 30 tokens at each step.
Inspired by the above observations, we propose Cascade Speculative Drafting (CS Drafting), a speculative execution algorithm that comprises multiple draft models, with the smallest being a statistical language model. Each neural draft model reviews generations from a smaller model and then proposes its reviewed content to either a larger draft model or the target model. In this design, the drafting of each neural model is accelerated by drafting from a smaller model, avoiding the inefficiency of autoregressive generation from neural models. We refer to this tiered speculative decoding approach as the Vertical Cascade. In addition, we suggest the use of smaller, faster draft models for generating high-rejection tokens that are trailing in drafting generation, forming the Horizontal Cascade. Along with the aforementioned Vertical Cascade, these strategies compose our complete CS Drafting approach, as illustrated in Figure 1.
@@ 2 Preliminary
The core concept of speculative decoding [14] involves the utilization of a small draft model for sequential token generation with validation by a larger target model resulting in reduced latency. This design accelerates sampling from autoregressive models without altering output distributions. At its heart, there are two key observations: 1) certain generations in language modeling are simpler than others and can be predicted by more efficient models correctly, and 2) using speculative execution along with a new sampling method enables faster, exact decoding from large models.
Specifically, let $x$ be the input tokens at a run and $\mathcal{M}_{t}$ and $\mathcal{M}_{d}$ are the target and the draft model respectively, $k$ be the number of draft tokens generated per step, and $\mathcal{M}_{t}(x)[i]$ and $\mathcal{M}_{d}(x)[i]$ be their probability output at $i$-th token when input is $x$. We interpret speculative sampling as a two-stage operation. In the proposing stage, we sample $\{x_{t+1},...,x_{t+k}\}$ from draft model $\mathcal{M}_{d}$ autoregressively and append them to $x$. In the reviewing stage, let $x_{i}\in\{x_{t+1},...,x_{t+k}\}$ represents the token at the current position, and we accept it if $\mathcal{M}_{d}(x)[i-1]\leq\mathcal{M}_{t}(x)[i-1]$; in the event that $\mathcal{M}_{d}(x)[i-1]>\mathcal{M}_{t}(x)[i-1]$, we reject $x_{i}$ with a probability of $1-\frac{\mathcal{M}_{t}(x)[i-1]}{\mathcal{M}_{d}(x)[i-1]}$ and proceed to resample $x_{i}$ from a recalibrated distribution $norm(\max(0,\mathcal{M}_{t}(x)[i-1]-\mathcal{M}_{d}(x)[i-1]))$ and reject any token following $x_{i}$. At the end, the target model will generate one additional token following the accepted tokens. Such a design guarantees the output is the same as sampling autoregressively using the target model alone [14].
@@ 3.1 Vertical Cascade
Additionally, we incorporate lenience, a hyperparameter that loosens the review process by the target model, allowing for faster speed at the trade-off of potentially differing results from the target model [14]. Lenience can be adopted during sampling or greedy decoding with speculative decoding. Let lenience $l\in[1,\infty)$. When sampling, the acceptance condition for token $x_{i}$ is transformed to $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$. If the acceptance condition is not satisfied, with a probability of $1-\frac{l\times\mathcal{M}_{t}(x)}{\mathcal{M}_{d}(x)}$, we reject $x_{i}$ and any following tokens. [^2] For simplicity, we assume the probability outputs are not standardized. We refer the readers to the speculative decoding paper [14] for the discussion on standardized sampling. When performing greedy decoding, the acceptance condition becomes deterministic and is simply either $argmax\mathcal{M}_{d}(x)[i]=argmax\mathcal{M}_{t}(x)[i]$ or $\mathcal{M}_{d}(x)[i]\leq l\times\mathcal{M}_{t}(x)[i]$.
@@ 4 Analysis
In this section, we provide theoretical analyses for Cascade Speculative Drafting. We begin with some notions. Let $\mathcal{M}_{t}$ be the target model, $\mathcal{M}_{d}$ be the draft model, and $k$ be the number of draft tokens generated per step.
@@
Despite the simple setting of EWIF, it is demonstrated that it aligns with the experimental results in most instances[14]. Therefore, our analysis will concentrate on EWIF.
@@ 4.1 Vertical Cascade
We analyze EWIF of vertical cascade using generating functions, a well-studied topic in combinatorial mathematics [24]. The properties of generating functions are useful in the recursion and evaluation process making our final expression simple.
@@
$\phi_{\alpha^{\prime},k}(\alpha)<\alpha$ for any $01>\alpha>0,1>\alpha^{\prime}>0,k>$ 0, so if $c_{d_{2}}\ll 1$, the EWIF of $\mathcal{M}_{d_{1}}$ and $\mathcal{M}_{d_{2}}$ is higher than EWIF of $\mathcal{M}_{d_{1}}$ alone.
@@
Therefore, with the statistical model having negligible cost (i.e., $c_{d_{2}}\ll 1$), it can almost always improve the efficiency of an SD system.
@@ 4.2 Horizontal Cascade
Using the information provided by Leviathan et al. [14], we calculate a simulated EWIF under the assumption that the event of acceptance by the target model is a Bernoulli trial. The results, shown in Table 1, indicate that speculative sampling with a horizontal cascade achieved better EWIF than vanilla speculative sampling under this assumption.
@@
Table 1: Simulated EWIF under the assumption that the acceptance distribution is a Bernoulli distribution. BASE and SMALL refer to FLAN-T5-base and FLAN-T5-small. In the simulation, speculative sampling with horizontal cascade exceeded the performance of the vanilla speculative decoding on both CNN Dailymail [16] and WMT EnDe [2] datasets.
@@ 5 Experiments
5.1 Experimental Setup
Metrics We use both our proposed standardized walltime improvement and walltime for evaluation:
  • Standardized walltime improvement (SWI) assumes each forward run of a model takes a constant amount of time which can be recorded data of previous work [14] or heuristics such as the total number of the parameters of a model. Under this assumption, the value of SWI is the speedup of the speculative method over autoregressive generation. SWI alleviates hardware variation and features full reproducibility of experiment results.
  • Walltime refers to the actual elapsed time taken to complete a specific task or operation in a real-world scenario. Despite being less reproducible and sensitive to noise, walltime better represents the performance for individual users. In our experiment, walltime is measured in the form of the number of tokens generated per second on our GPU.
Dataset Algorithm $\{\mathcal{M}_{d_{i}}\}$ Speedup (MS) Speedup (PW)
GSM8K Autoregressive - 1 1
[…]
MMLU Autoregressive - 1 1
[…]
@@ 5.1 Experimental Setup
Table 2: The experimental results on FLAN-T5. Speedup (MS) is the standardized walltime improvement with the assumption that the latency of each run of a model is its number of parameters (model size). Speedup (PW) is the SWI with the assumption that the latency of each run of a model is the time cost data reported from previous work [14].
@@
Datasets We chose two commonly used datasets for our experiments. For both datasets, we conducted experiments in a zero-shot chain-of-thought setup [13, 23]:
@@
  • GSM8K [7] is a dataset comprising 8,500 high-quality, linguistically diverse, grade-school math word problems. It focuses on multi-step reasoning with problems that are typically solvable using basic arithmetic in 2 to 8 steps.
@@
  • MMLU [10], or Massive Multitask Language Understanding, is a benchmark for testing how well large language models grasp knowledge. It encompasses 57 diverse subjects, ranging from elementary science to advanced law.
Baselines To verify the effectiveness of both vertical and horizontal cascade strategies intuitively, we first compare the performance of CS Drafting with different numbers of cascades, as well as its performance against standard speculative decoding [14]. Additionally, CS Drafting can also operate vertically by combining with other advanced decoding methods. To verify this, we also leverage tree attention, as used in Medusa [3], and compare its performance with Medusa.
Implementation Details To ensure the generality of our findings, we perform experiments on both encoder-decoder and decoder-only models. For encoder-decoder models, we choose our target and draft models from the FLAN-T5 [6] family for our experiment, as there is a large variation in model sizes within the FLAN-T5 family (ranging from 77 million to 11 billion parameters). We use FLAN-T5-xxl as our target model, FLAN-T5-base and FLAN-T5-small as our reviewing draft models. For decoder-only models, we select Vicuna-7B [28], a fine-tuned version of LLaMA [20] as the target model. We use a 68M model with the same tokenizer as the reviewing draft model [^3] https://huggingface.co/double7/vicuna-68m. We also leverage tree attention [15, 3] with CS Drafting for the experiments on Vicuna-7B. In both cases, the Max-Gram algorithm is used as the generating draft model. Since we do not observe any significant difference between sampling with temperature $1$ and greedy decoding in previous speculative decoding experiments [14], and to ensure our experiments are fully reproducible, we perform sampling at temperature 00, i.e., using greedy decoding by default. To align our experiment with current common usage, we do not perform fine-tuning for CS Drafting, and the generation is conducted in a zero-shot manner. We include hyperparameter details in Appendix B. All of our experiments involving walltime are performed on a single NVIDIA A40 GPU.